OpenStudy (anonymous):

Simplify. Remember to use absolute - value notation when necessary. If a root cannot be simplified, state this. -3 √ -1000

OpenStudy (anonymous):

√ -a = √ (-1*a)=√ a√ -1=√ a*i

OpenStudy (anonymous):

I have -30i √ 10 and i am told that You can multiply something by itself 3 times and get a negative, so this does not result in an i for the radical.

OpenStudy (anonymous):

I am lost with what is being asked

OpenStudy (anonymous):

hmm... i^2 = -1 -1000 = -10^3 but the root is square, so I don't see how this could help